Solve the system by addition or substitution.
3x – 4y = 8
6x – 2y = 10

Found 2 solutions by checkley71, praseenakos@yahoo.com:
Answer by checkley71(8403)

3X-4Y=8 MULTIPLY THIS EQUATION BY -2 & THEN ADD
-6X+8Y=-16
6X-2Y=10
------------------------------------
6Y=-6
Y=6/-6
Y=-1 SOLUTION
3X-4*-6=8
3X+24=8
3X=8-24
3X=-16
X=-16/3 SOLUTION

Answer by praseenakos@yahoo.com(507)

QUESTION:
Solve the system by addition or substitution.
3x – 4y = 8
6x – 2y = 10

ANSWER:
.
3x – 4y = 8 -------------------------(1)
6x – 2y = 10 --------------------------(2)
Multiply first equation 2
==> 2 ( 3X - 4Y ) = 2 * 8
==> 2*3x - 2*4y = 16
==> 6x - 8y = 16 ------------------------(3)

Subtract equation (2) from equation (3)
==> 6x - 8y -( 6x - 2y )= 16 -10
==> 6x - 8y - 6x + 2y = 6
==> 6x - 6x - 8y + 2y = 6
==> -6y = 6
==> y = 6 /-6
==> y = -1


Now input the value y = -1 in first equation.
==> 3x – 4 *-1 = 8
==> 3x + 4 = 8
==> 3x = 8-4
==> 3x = 4
==> x = 4/3


So the solution is,
x= 3/4 and y = -1

Hope you understood.

Regards.
praseenakos@yahoo.co.in